OpenStudy (anonymous):

does (H2 + B26)^4 = (H2^4 + B26^4) or H2B26 ^4?

OpenStudy (anonymous):

H2 and B26 are just variables, i guess so. than for easier handling let H2=x and B26=y (x+y)^4=x^4+4x^3y+6x^2y^2+4xy^3+y^3 so your answer cannot be correct